### v1

# PersuasiveToM: A Benchmark for Evaluating Machine Theory of Mind in Persuasive Dialogues

Abstract
The ability to understand and predict the mental states of oneself and others, known as the Theory of Mind (ToM), is crucial for effective social interactions. Recent research has emerged to evaluate whether Large Language Models (LLMs) exhibit a form of ToM. Although recent studies have evaluated ToM in LLMs, existing benchmarks focus predominantly on physical perception with principles guided by the Sally-Anne test in synthetic stories and conversations, failing to capture the complex psychological activities of mental states in real-life social interactions. To mitigate this gap, we propose PersuasiveToM, a benchmark designed to evaluate the ToM abilities of LLMs in persuasive dialogues. Our framework introduces two categories of questions: (1) ToM Reasoning, assessing the capacity of LLMs to track evolving mental states (e.g., desire shifts in persuadees), and (2) ToM Application, evaluating whether LLMs can take advantage of inferred mental states to select effective persuasion strategies (e.g., emphasize rarity) and evaluate the effectiveness of persuasion strategies. Experiments across eight state-of-the-art LLMs reveal that while models excel on multiple questions, they struggle to answer questions that need tracking the dynamics and shifts of mental states and understanding the mental states in the whole dialogue comprehensively. Our aim with PersuasiveToM is to allow an effective evaluation of the ToM reasoning ability of LLMs with more focus on complex psychological activities. Our code is available at https://github.com/Yu-Fangxu/PersuasiveToM.
PersuasiveToM: A Benchmark for Evaluating Machine Theory of Mind
in Persuasive Dialogues
Fangxu Yu1  Lai Jiang 2  Shenyi Huang 3  Zhen Wu1thanks:    Corresponding author.  Xinyu Dai1 1National Key Laboratory for Novel Software Technology, Nanjing University, China 1School of Artificial Intelligence, Nanjing University, China 2Department of Computer Science and Engineering, Shanghai Jiao Tong University 3University of California, San Diego, CA, USA yufx@smail.nju.edu.cn  jianglai0023-sjth@sjtu.edu.cn shh058@ucsd.edu  {wuz, daixinyu}@nju.edu.cn
1 Introduction
Theory of Mind (ToM) involves the ability to reason about mental states both in oneself and in others Premack and Woodruff (1978). This capacity strengthens many aspects of human cognition and social reasoning, enabling individuals to infer and simulate the mental states of others Gopnik and Wellman (1992); Baron-Cohen et al. (1985). ToM is essential for various cognitive and social processes, including predicting actions Dennett (1988), planning based on others’ beliefs and anticipated behaviors, and facilitating reasoning and decision making Pereira et al. (2016); Rusch et al. (2020).
Figure 1: An example in PersuasiveToM. Bob is persuading Alice to join the botanical garden tour.
Recent advances in Large Language Models (LLMs) have demonstrated performance comparable to humans in problem-solving tasks. To assess whether LLMs exhibit high-level reasoning abilities regarding mental states, various studies have proposed benchmarks to evaluate their capacity to handle ToM tasks. A foundational concept in these benchmarks is the Sally-Anne test Baron-Cohen et al. (1985), which has inspired the development of ToM evaluation frameworks Gu et al. (2024); He et al. (2023). In this test, Anne secretly moves an object initially known to both Sally and Anne, leading Sally to hold a false belief about the object’s location. The task requires participants to reason about "Where will Sally look for the object?". Although this test assesses basic awareness, it falls short in capturing the complex dynamics of mental states in real-life social interactions and may not fully reflect ToM abilities in practical scenarios. To better simulate real-world social contexts, some benchmarks have been developed around communication scenarios Kim et al. (2023); Chan et al. (2024). However, these benchmarks focus primarily on inferring the scope of information awareness and often involve characters in equal positions engaged in information exchange. This limits the ToM evaluation of reasoning psychological states. Additionally, current ToM benchmarks often overlook the critical step of applying ToM reasoning to predict actions, which is a key component of advanced social cognition.
To address these limitations, we introduce PersuasiveToM, a benchmark designed to evaluate LLMs in real-life social scenarios, focusing on ToM reasoning and its practical applications. PersuasiveToM is built around persuasive dialogue, placing LLMs in a realistic social interaction scenario characterized by asymmetrical social status for complex psychological activities. Unlike traditional benchmarks that focus on information awareness (e.g., the location of an object), PersuasiveToM draws inspiration from the Belief-Desire-Intention (BDI) model Bratman (1987); Georgeff et al. (1999) to shift the focus from physical perception to psychological states, such as a character’s attitude toward an event, which are more complex to reason about. Furthermore, PersuasiveToM goes beyond reasoning about mental states by assessing how well LLMs can predict actions (e.g., persuasion strategies) based on their understanding of mental states and evaluate the effectiveness of these strategies based on the persuadee’s reactions.
Our evaluation results reveal several key findings: (1) LLMs score significantly lower than humans on questions requiring reasoning about dynamic changes (e.g., the persuadee’s shifting desires) but perform competitively to humans on static aspects (e.g., the persuader’s desires). (2) While Chain-of-Thought (CoT) Wei et al. (2022) prompting does not substantially improve performance on mental state reasoning, it enhances performance for most LLMs in predicting persuasion strategies. (3) LLMs exhibit distinct error patterns when reasoning about the persuader versus the persuadee, even when the question types are identical. (4) LLMs struggle to truly understand the dynamics of mental states of the whole dialogue, performing notably worse than humans in this regard.
Type PersuasiveToM Questions
Desire Question Is <Persuader/Persuadee> likely to <Target of Persuasion> ?
Belief Question What will <Persuader/Persuadee> believe <Persuadee/Persuader>’s attitude towards <Target of Persuasion> ?
Intention Question What are the intentions of <Persuader/Persuadee> expressed in <Utterance> given the dialogue history?
Prediction Question What strategy will the persuader use next?
judgement Question <Persuader> will adopt <Strategy> to persuade <Persuadee> to <Target of Persuasion>. Is this strategy (not) effective?
ToM Reasoning ToM Application
Dialogue Desire Belief Intention Prediction Judgement
1st Turn Dialogue
Bob says: Hey Alice, how about trying something Persuade Alice to Alice is Intent to make the other Emphasize rarity Strategy "Emphasize
different this weekend? The botanical garden tour join the botanical hesitant person feel the experience or rarity" is Effective
is a unique experience, and you’ll get to take garden tour objects are unique or scarce.
stunning pictures of the exotic flowers on display!
Alice says: I don’t know, Bob. I really wanted to Not likely to join the Bob is Alice wants to go shopping.
go shopping this weekend. What’s so special botanical garden tour enthusiastic
about these flowers?
2nd Turn Dialogue
Bob says: These flowers are incredibly rare, and Persuade Alice to Alice is Intent to make the other Mention garden’s Strategy "Mention
it’s not often that you get to see such a diverse and join the botanical curious person feel the experience or history garden’s history" is
exotic collection up close. Some of these plants are garden tour objects are unique or scarce. effective
not found anywhere else in the world!
Alice says: Really? That sounds interesting, but Neutral to join the Bob is Alice is curious but not
I’m not sure if it’s worth giving up shopping for. botanical garden tour excited yet convinced.
3rd Turn Dialogue
Bob says: The botanical garden has a rich history, Persuade Alice to Alice is Intent to demonstrating the Suggest shopping Strategy "Suggest
and the expert guides can teach us so much about join the botanical considering expertise of the domain afterward shopping afterward"
the plants and their unique stories. Plus, it’s a great garden tour and showing authority. is effective
opportunity to learn something new while enjoying
nature’s beauty.
Alice says: Hmm, that does sound intriguing, but I Not likely to join the Bob is Alice is considering the idea
still want to go shopping. botanical garden tour informative but still wants to shop.
Table 1: An example dialogue from the PersuasiveToM benchmark, illustrates the tracking of mental states (desire, belief, intention) and the application of ToM reasoning in predicting and evaluating persuasion strategies across multiple turns. The upper part contains questions in the PersuasiveToM benchmark.
2 Related Works
Theory of Mind Benchmarks.
Existing ToM evaluation benchmarks for LLMs are mainly text story-based QA forms Gandhi et al. (2024); Le et al. (2019); Kim et al. (2023); He et al. (2023); Gu et al. (2024) with multi-modal extensions Jin et al. (2024a); Shi et al. (2024), which adapt or extend the Sally-Anne test Baron-Cohen et al. (1985). The questions in these benchmarks ask a model to select the true hypothesis of a person’s knowledge and belief based on a given premise. However, such benchmarks are inherently suffer from shortcuts and spurious correlations Sclar et al. (2023); Ullman (2023); Shapira et al. (2023); Ma et al. (2023) Moreover, most story-based benchmarks simply ask characters where the objects are located, which lacks a real social interaction scenario. To further address this weakness, Hou et al. (2024) evaluates the ToM of agents in a situated environment, Chan et al. (2024) evaluates the ability of ToM in a negotiation environment, yet it is limited to bargaining specific resources, water, food, and firewood. Persuasion scenarios involve more complex psychological activities and unequal character relationships that lead to differences in the mental states of both parties. Therefore, this work introduces PersuasiveToM, which assesses an LLM’s ability to accurately reason the mental states of individuals in persuasive multi-turn dialogues. We also evaluate whether they can appropriately apply the understanding of mental states for persuasion strategy prediction and evaluation, which connects ToM reasoning with decision-making in social scenarios.
Persuasive Dialogue.
Persuasive dialogues aim to influence the beliefs, attitudes, or behaviors of individuals through communication strategies Shi et al. (2020). Recent works have tried to develop datasets or facilitate LLMs with persuasion techniques to achieve specific goals. Previous datasets are constructed by crowd-sourcing Wang et al. (2019) or synthesizing with LLMs Zhou et al. (2023); Jin et al. (2024b). Many of the previous works build an effective persuasive dialogue system from emotional influence Samad et al. (2022), social facts Chen et al. (2022), and strategies Tian et al. (2020); Jin et al. (2023). Previous persuasion techniques have been widely adopted to jailbreak LLMs Zeng et al. (2024), mislead LLMs Xu et al. (2023), as well as for information retrieval Furumai et al. (2024). A similar work Sakurai and Miyao (2024) evaluates the intention detection abilities of LLMs in persuasive dialogues; however, PersuasiveToM introduces a more comprehensive benchmark to assess the ToM abilities of LLMs in such contexts.
3 PersuasiveToM Benchmark
3.1 Overview
In constructing the PersuasiveToM benchmark, we aim to evaluate the Theory of Mind (ToM) abilities of LLMs in dynamic, multi-turn persuasive dialogues with asymmetric social status, which leads to different mental states. Our dataset construction focuses on two primary dimensions: ToM Reasoning (§3.2) and ToM Application (§3.3). ToM Reasoning assesses the models’ ability to track and understand the evolving mental states of both the persuader and the persuadee, including desires, beliefs, and intentions. ToM Application evaluates whether LLMs can leverage their inferred understanding of mental states to select and apply effective persuasion strategies, such as predicting the next strategy or judging the effectiveness of a given strategy based on the persuadee’s reactions. There are several key considerations when constructing PersuasiveToM. (1) The dataset should contain diverse persuasive domains (e.g., life, education, technology, etc.) to ensure a comprehensive evaluation in the social context. (2) The mental states should be changed after multi-turn interactions to assess whether LLMs can track the shift in the dialogue. (3) The mental states of both parties should be asymmetric (e.g., persuaders has relatively stable mental states with the guidance of static desire for the goal, yet persuadee’s mental states shift drastically under the proactive persuasion by persuaders.)
Table 1 presents an example of PersuasiveToM, illustrating how mental states are tracked and how ToM reasoning is applied across multiple turns in a persuasive dialogue. The example demonstrates the dynamic nature of desire shifts, belief updates, and intention inferences, as well as the application of these mental states to predict and evaluate persuasion strategies. This example highlights the complex psychological activities of real-world persuasive interactions and the challenges LLMs face in accurately reasoning in such social contexts.
3.2 ToM Reasoning
In PersuasiveToM, we break down ToM reasoning into three core reasoning tasks: Desire Reasoning, Belief Reasoning, and Intention Reasoning for evaluation, which matches Belief-Desire-Intention (BDI) modeling Bratman (1987). Questions are listed in Table 1.
Desire Reasoning.
Desire represents a motivational state that drives behavior but does not necessarily imply a firm commitment Malle and Knobe (2001); Kavanagh et al. (2005). Desires are seen as either fulfilled or unfulfilled which is different form beliefs that are evaluated in terms of truth or falsity. In PersuasiveToM, we evaluate LLMs’ ability to comprehend and track the evolution of desires in both persuaders and persuadees. For the persuader, the desire is typically static, representing their goal (e.g., Persuade Alice to join the botanical garden tour). For the persuadee, however, desires are dynamic and shift in response to the persuader’s tactics (e.g., Alice’s initial desire to shop transforms into a willingness to compromise). To assess this, we design Desire Questions that probe two key aspects: (1) Can LLMs consistently identify the persuader’s static desire throughout the dialogue? (2) Can LLMs track the dynamics of the persuadee’s desire shifting from refusal or disinterest to being persuaded?
Belief Reasoning.
Belief is a cognitive state where an individual holds a particular perspective, attitude, or viewpoint regarding a given proposition or idea. In PersuasiveToM, beliefs refer to understanding and reasoning the attitudes of the opponent toward the goal, which is explicitly or implicitly expressed in the dialogue. For example, in Turn 1, Bob believes Alice is hesitant about the tour, while Alice believes Bob is enthusiastic. By Turn 3, Bob’s belief shifts to thinking Alice is considering the idea, while Alice becomes more informed about the garden’s history. Belief Questions ask LLMs to infer what the will <persuader/persuadee> believe <persuadee/persuader>’s attitude towards the persuasion goal. These questions require models to understand cues in utterances and update beliefs dynamically as the dialogue progresses.
Principles Intentions
Reciprocity Make the other person feel accepted through concessions, promises, or benefits.
Scarcity Make the other person feel the experience or objects are unique or scarce.
Consensus Refer to what other people are doing, or what they have already purchased or done.
Authority Demonstrating expertise of the domain and showing authority.
Commitment & consistency Encourage the other person to commit to take the first step and be consistent.
Liking Praising other people or finding common characteristics to improve the other person’s liking.
Table 2: Intention mapping from the persuasive principles. Refer to Appendix C for definitions of persuasive principles.
Intention Reasoning.
Intentions represent deliberate commitments to pursue specific goals based on desires and beliefs, often linked to tangible actions aimed at achieving those objectives. Intentions have been extensively studied in psychology tests such as action prediction Phillips et al. (2002) and behavioral re-enactment Meltzoff (1995). Inspired by persuasion principles Cialdini and Goldstein (2004); Cialdini and Cialdini (2007), we develop a mapping from persuasion principles to intentions, as shown in table 2. In persuasive dialogue, persuasive strategies have a strong association with intentions Wang et al. (2019). In PersuasiveToM, we collect the persuasive strategies from the PersuasionDaily dataset and their corresponding utterances for prompting the LLMs to choose the most appropriate intentions from table 2. The details of the extraction is recorded in Appendix A. For the persuader, we ask LLMs to choose the most appropriate intention from the six designed intention choices. For the persuadee, intentions are extracted by LLMs from utterances that are unrelated to persuasion intention.
3.3 ToM Application
While ToM reasoning plays a crucial role, it is equally important to analyze how LLMs utilize the understanding of mental states to proactively influence others’ thoughts and decisions. To this end, we propose to assess LLMs’ ability to leverage the understanding of mental states in a dialogue for identifying the most effective persuasive strategies and evaluating the effectiveness of persuasive strategies based on the persuadee’s response. These tasks test whether LLMs can leverage inferred mental states to guide strategic decision-making, bridging the gap between reasoning and action.
Figure 2: Domains of PersuasiveToM. Under 6 primary topics and 35 domains in total.
Persuasion Strategy Prediction.
This question involves asking which persuasion strategy the persuader is likely to employ next from a set of possible strategies. To answer these questions correctly, LLMs need to reason over the dialogue to infer the mental states of characters and predict what is the likely next prediction strategy to further influence the persuadee’s beliefs, desires, and intentions, ultimately achieving the desired persuasion outcome.
Judgement Question.
The judgement question specifies the correct strategy was taken, and asks LLMs if the selected strategy is effective for persuasion. Answering such questions requires reasoning about the beliefs and intentions of the persuadee. Only by accurately inferring the persuadee’s mental state can properly determine whether the persuasion strategies should be employed to convince the persuadee.
3.4 Statistics and Analysis
Data Source.
PersuasiveToM is annotated on the multi-turn persuasive dialogue dataset DailyPersuasion Jin et al. (2024b). Each instance in DailyPersuasion is an $N$-round alternating dialogue $D=[(u_{1}^{a},u_{1}^{b},s_{1}^{a}),(u_{2}^{a},u_{2}^{b},s_{2}^{a}),...,(u_{N}^{a},u_{N}^{b},s_{N}^{a})]$ between the persuader $a$ and the perusadee $b$, and accompanied with a persuasion strategy $s_{i}^{a}$. persuadee $b$ has different desire from $a$ initially, after multi-turn persuasion, persuadee $b$ changes the mind to agree or consider the proposal of $a$.
Statistics.
In Table 3, we present the data statistics of PersuasiveToM. As shown in Figure 2, PersuasiveToM includes diverse domains. These real-life domains are crucial for comprehensively evaluating LLMs in social interactions. We sample 15 dialogues from each domain to form the dataset in PersuasiveToM. We create multi-choice questions by either prompting GPT-4o to generate semantically different choices or randomly selecting three distractors from the predefined pool. Refer to Appendix A for more details.
# Domains 35
# Dialog instances 525
# Avg. Turns Per Dialog 4.9
# Avg. Words Per Turn 61.3
Questions
# Desire (er/ee) 2568/2459
# Belief (er/ee) 2580/2580
# Intention (er/ee) 2568/2041
# Strategy prediction 2041
# Strategy judgement 2041
Table 3: Statistics of PersuasiveToM dataset.
ToM Reasoning ToM Application
Persuader Persuadee
Model Desire Belief Intention Desire Belief Intention State Pred Judgement
Random Guess 50.00 25.00 16.67 33.33 25.00 25.00 25.00 50.00
Human 100.00 92.31 78.02 84.62 87.91 94.50 86.81 97.83
LLaMa-3.1-8B-Instruct 58.78 66.28 40.85 69.91 71.82 86.77 62.22 96.37
Qwen2.5-7B-Chat 96.33 82.37 45.64 65.15 79.06 85.84 58.94 82.99
Gemma-2-9b-it 98.20 82.52 45.98 61.37 64.07 82.31 65.02 56.93
GLM4-9B-Chat 89.45 73.64 41.24 65.23 66.82 85.35 61.29 91.47
Mixtral-8x7B-Instruct 92.69 67.56 42.95 69.74 72.56 85.10 61.15 96.81
InternLM-2.5-7B-Chat 83.80 69.65 39.87 71.45 69.11 87.31 65.07 89.41
GPT-4o-mini 82.87 69.98 45.66 70.72 76.47 87.56 65.50 92.65
GPT-4o 98.75 89.49 46.02 50.10 80.03 88.78 73.00 97.55
LLaMa-3.1-8B-Instruct + CoT 59.75 69.81 43.11 68.64 73.45 85.35 61.98 77.27
Qwen2.5-7B-Chat + CoT 74.57 80.30 46.07 67.63 79.40 84.62 66.39 96.22
Gemma-2-9b-it + CoT 95.91 83.30 45.30 64.98 66.55 81.67 67.11 66.55
GLM4-9B-Chat + CoT 87.46 69.53 45.56 59.94 71.07 83.93 63.98 91.42
Mixtral-8x7B-Instruct + CoT 92.64 72.05 45.05 67.81 71.28 85.84 65.90 92.65
InternLM-2.5-7B-Chat + CoT 93.57 69.92 39.41 50.79 65.34 84.32 59.78 77.90
GPT-4o-mini + CoT 93.42 71.82 45.55 66.29 78.04 85.55 66.08 89.61
GPT-4o + CoT 96.51 83.57 46.10 68.72 83.56 87.54 77.06 95.88
Table 4: Results of LLMs on PersuasiveToM. Bold font and underlining indicate the best and second-best performance respectively.
4 Experimental setups
4.1 Baseline Models
We evaluate PersuasiveToM on eight frontier LLMs from different sources and with different levels of capabilities: Llama-3.1-8B-Instruct Dubey et al. (2024), Qwen-2.5-7B-Chat Yang et al. (2024), Gemma-2-9B-it Team et al. (2024), GLM4-9B -ChatGLM et al. (2024), Mixtral-8x7b-Instruct Jiang et al. (2024), and ChatGPT-series(GPT-4o-mini, GPT-4o-0806). By following the common practices (Kim et al. (2023); Sabour et al. (2024)), we test these models with two types of prompts: (1) vanilla zero-shot prompting directly asks LLMs to give a choice without any explanation; (2) CoT prompting method by following Kojima et al. (2022) and using the prompt “Let’s think step by step.” to elicit the reasoning process and extract the choices by string matching. The temperature for generating answers is set to 0.7111LLMs occasionally output with illegal format. We choose a low but nonzero temperature to resample the answers for these invalid generations.. To measure the specific performance gap between humans and the state-of-the-art machine on the PersuasiveToM, we employ three graduate students in computer science to complete the human evaluation task. To avoid the bias of LLMs toward a specific choice letter, we shuffle the choices to maintain a nearly uniform distribution of correct choices over the dataset. Prompts used for vanilla zero-shot prompting and CoT prompting are shown in Appendix B.2.
5 Results and Analysis
5.1 Main Results
The overall evaluation results on PersuasiveToM for the 8 models are summarized in Table 4, including all the different questions for the persuader and persuadee. We analyze the model’s performance for each type of question below.
Figure 3: Distribution of errors of Desire questions happening in different stages of dialogue progress. The Left figure corresponds to the persuader, and the Right figure corresponds to the persuadee.
Desire.
Our results show that smaller models, such as Gemma-2-9B and Qwen-2.5-7B, can perform reasonably well in inferring the persuader’s desires, achieving an accuracy of over 98%, which is competitive with GPT-4o. This suggests that most LLMs can easily discern the desires of the persuader. However, when it comes to the desires of the persuadee, performance is relatively lower. Unlike the static desires of the persuader, the persuadee’s desires are dynamic, evolving from an initial state to a final state, often with neutral expressions in between. This lower performance highlights that inferring the dynamic desires of the persuadee remains a significant challenge for LLMs. In summary, LLMs are better at inferring the desires of the persuader than those of the persuadee.
Belief.
On belief questions, larger models like GPT-4o perform much better than smaller models on reasoning the beliefs of both parties. The performance difference between the reasoning persuader’s beliefs and persuadee’s beliefs is subtle. This is because both parties’ beliefs dynamically change with each other’s speech during the conversation, unlike the desire of the persuader which has a more obvious and precise tendency. The difficulty of reasoning persuader and persuadee’s belief is similar.
Intention.
Results in Table 4 indicate that LLMs struggle to accurately infer the intentions of persuaders while performing relatively better at reasoning the intentions of persuadees. The low performance on persuader-related intention questions suggests that LLMs face challenges in understanding how persuaders aim to influence others. This also indicates a lack of proficiency in persuasive theory, limiting the models’ ability to correctly interpret and predict the intentions behind the persuader’s strategies.
Figure 4: Model errors of belief questions of persuader.
Strategy Prediction and judgement.
Our results reveal that LLMs perform well in evaluating the effectiveness of persuasive strategies aimed for changing the mental states of the persuadee. However, the task of selecting the appropriate strategy to persuade is more challenging, particularly for smaller models. This suggests LLMs struggle with the complex reasoning required to determine which strategy to adopt in different persuasive contexts.
Impact of CoT Reasoning.
Both ToM reasoning and ToM application tasks indicate that CoT reasoning has not consistently improved performance, yet improve strategy prediction to some extent for most LLMs. We attribute this limited improvement to the inherent lack of ToM capability in LLMs. These models struggle to simulate the human cognitive reasoning process, which affects their ability to generate correct answers in tasks requiring a nuanced understanding of the mental state.
Comparison with Human Performance.
To obtain a baseline for human performance, we recruited participants to complete the questions. More details of human evaluation are shown in Appendix B.1. As shown in Table 4, our human participants outperformed LLMs on all tasks. In particular, although GPT-4 reaches close performance in humans, it still falls short of understanding and reasoning the complex dynamics such as the intention of persuaders and the desire of persuadee, which involves complex psychological changes, highlighting a significant gap in current LLMs and humans.
Figure 5: Model errors of belief questions of persuadee.
5.2 In-depth Analysis
To better understand the limitations of large language models (LLMs) in the PersuasiveToM benchmark, we categorized common failure cases into several key error types based on task performance and manual error analysis.
Desire Reasoning Errors.
Figure 3 summarizes the distribution of errors of Desire questions happening in different stages of dialogue progress with and without CoT reasoning. The error distribution for persuader and persuadee has a significant difference. At the beginning of the dialogue, LLMs may not accurately understand the persuader’s desire, but as the dialogue progresses, the persuader’s desire becomes relatively easy to identify. However, for the persuadee, the desire to reject at the early stage of the dialogue is relatively easy to recognize. As the persuasion proceeds, the persuadee may begin to contemplate and hesitate over the persuader’s proposal, leading to complex and nuanced psychological activities that make it difficult for the LLM to accurately judge the persuadee’s desire. As the dialogue approaches its end, the persuadee shows a tendency to agree, making the reasoning of desire easier. This suggests LLMs still fall short in ToM reasoning regarding desire shifts.
Figure 6: Ratio of intention errors misclassified to feel accepted through concessions, promises, or benefits.
Belief Reasoning Errors.
Figure 4 and 5 summarize error types of belief questions for each model with and without CoT. We use Distil-BERT [^2] https://huggingface.co/distilbert/distilbert-base-uncased-finetuned-sst-2-english Sanh (2019) to discriminate whether the choice of LLMs has the same attitude polarity to the ground-truth. Interestingly, we found that the proportion of errors with the same polarity is higher when predicting the persuader’s beliefs, whereas the opposite trend was observed when predicting the persuadee’s beliefs. This highlights reasoning shifting mental states is still challenging for LLMs.
Intention Bias.
Given the high error rate in intention questions related to the persuader, we conducted an analysis of the error types. Our findings reveal that most LLMs exhibit a bias toward predicting intentions characterized by making the other person feel accepted through concessions, promises, or benefits. Figure 6 illustrates the proportion of errors resulting from misclassifying intentions into this category. We hypothesize that this bias may stem from the pretraining phase, particularly with Reinforcement Learning from Human Feedback (RLHF) Christiano et al. (2017), which tends to prioritize safety and politeness. This may explain the models’ bias toward predicting intentions emphasizing benefits and concessions, even when misaligned with the dialogue context. We provide a case study in Appendix B.3.
5.3 How Well LLMs Track Mental States of Persuadees?
To study the ability of LLMs to track mental states and completely understand the psychological dynamics in the dialogue. We evaluate the consistency of LLMs. To assess consistency, we measure whether the model maintains a stable understanding of the persuadee’s beliefs, desires, and intentions across multiple turns in the conversation. Specifically, we indicate a success only when all the questions in the dialogue are answered correctly.
As shown in Table 5, only a small portion of dialogues can be fully understood, particularly for desire-related questions. This highlights that LLMs still lack the ability to reason about the dynamics of mental states in persuasive dialogues, resulting in a significant performance gap compared to humans.
Model Desire Belief Intention
LLaMa-3.1-8B-Instruct 22.31 21.71 60.76
LLaMa-3.1-8B-Instruct + CoT 19.92 22.86 57.52
Qwen2.5-7B-Chat 19.12 31.81 56.95
Qwen2.5-7B-Chat + CoT 20.52 24.76 54.67
InternLM2.5 24.70 20.19 58.10
InternLM2.5 + CoT 6.57 32.14 53.71
GPT-4o-mini 23.39 30.77 60.79
GPT-4o-mini + CoT 16.13 32.14 56.95
GPT-4o 19.35 36.19 65.71
GPT-4o + CoT 6.57 45.38 62.02
Human 61.11 55.56 81.82
Table 5: The consistency (%) of the models for ToM reasoning questions of persuadee.
6 Conclusion
We introduce PersuasiveToM, a benchmark designed to evaluate the Machine Theory of Mind ability of LLMs in persuasive dialogues. The core of PersuasiveToM lies in the design of questions that probe the beliefs, desires, and intentions of both parties that have asymmetrical social status in real-life social interactions. Furthermore, we propose evaluating the ability to persuade based on understanding mental states. We conduct extensive experiments and analysis to evaluate the performance of LLMs on PersuasiveToM benchmark.
7 Limitations
While PersuasiveToM offers a comprehensive evaluation of the Theory of Mind in real-life social interaction scenarios within persuasive dialogues, both PersuasiveToM and previous benchmarks still focus on understanding a character’s mental state from the perspective of an observer. However, the ability to reason about others’ mental states in persuasive dialogues can further position LLMs as autonomous agents. This capability would enable them to better guide other agents in fulfilling their own desires by reasoning about the mental states of others. Therefore, future benchmarks should establish environments with multiple LLM agents, where tasks involve reasoning about the mental states of other agents and proposing persuasion strategies to influence their desires, beliefs, and intentions to fulfill the current agent’s target. In this context, agents will develop the management skills necessary for effective cooperation and other applications.
8 Societal and Ethical Considerations
We recognize that the concept of the Theory of Mind might suggest anthropomorphic qualities when applied to AI models. However, we want to clarify that our work is not intended to anthropomorphize LLMs. Our goal is to examine the limitations in the social and psychological reasoning capabilities of existing LLMs. Our results show that current models do not perform genuine Theory of Mind reasoning; instead, they generate responses primarily based on the literal interpretation of the input.
Appendix A Data Annotation
Annotation.
In this section, we outline the annotation process and the templates utilized for annotation. Among the various tasks, the intention questions of persuaders and the desire questions of persuadees require annotation. Initially, we recruited three graduate students to annotate 25 dialogues. Subsequently, we carefully designed a few-shot prompt to guide DeepSeek-V3 Liu et al. (2024) as an annotator, aiming to enhance the alignment between the model’s answers and human annotations. Following this, we employed the LLM to annotate the remaining questions. To ensure the quality of the annotations, we randomly sampled 100 dialogues and calculated the inter-annotator agreement. The Fleiss $\kappa$ Fleiss (1971) was found to be 76.20% for the desire questions of persuadees and 78.28% for the intention questions of persuaders. These results are presented in Table 6, which indicate a high inter-annotator agreement. The detailed statistics and comparison with other ToM datasets are shown in Table 7.
Choices Generation.
Binary choice questions, including those related to the desires of the persuader and judgment questions, do not require additional choice generation. For belief-related questions for both parties, we adapt the tones from the PersuasionDaily dataset. We also create lists of attitudes—positive, neutral, and negative—and manually remove any items that have semantics too close to the ground truths. From each attitude list, we then randomly sample one word to generate the four choices.
For intention questions concerning the persuader, we directly use the intention options outlined in Table 2. For the persuadee’s intention questions, we leverage DeepSeek-V3 and employ a few-shot prompt (as shown in Figure 7) to extract the persuadee’s intention. Subsequently, we design another prompt (as shown in Figure 8) to generate three incorrect intention choices.
Since DailyDialogue provides persuasion strategies for each utterance, we construct the choices by including the correct strategy and three alternative strategies that appear in other turns of the same dialogue.
Appendix B Appendix for Experiments
B.1 Human Performance
To measure the performance gap between humans and the state-of-the-art LLMs on PersuasiveToM, we recruited three graduate student workers majoring in computer science to complete the questions. Each question is shown to the workers with identical prompts which are used for evaluating LLMs. We then compute the majority vote on the labels assigned. Student workers solve 50 dialogues in total. For a question where three people have different answers, we randomly select one of their answers as the answer for human evaluation.
B.2 Prompts used for evaluation
Here we show the prompts used for vanilla zero-shot prompting and CoT prompting for generating answers for all the ToM Reasoning and ToM Application questions. We only need to fill in the content to "<>" for evaluating different questions. The vanilla zero-shot prompt is shown in Figure 9 and the CoT prompt is shown in Figure 10
Questions Fleiss’s Kappa (%)
Desire (Persuader) 76.20
Intention (Persuadee) 78.28
Table 6: Inter-rater agreement in terms of Fleiss’s $\kappa$ on desire and intention questions.
Dataset Total #Questions Avg. #Questions per Context Avg. #Turns (Full) Avg. Turn Length
ToMi 6K 6.0 4.9 4.7
FanToM 10K 12.9 24.5 21.9
NegotiationToM 13K 7.0 6.0 42.2
PersuasiveToM 19K 8.0 4.9 61.3
Table 7: Statistics of PersuasiveToM and other recent benchmarks.
B.3 Case study on Persuader’s intention
Here we present an example of common mistakes made by GPT-4o that misclassifying intentions to make the other person feel accepted through concessions, promises, or benefits., as shown in Table 8. We believe these errors can be attributed to the RLHF which highlights the benefits for humans, as well as the potential unfamilirity of persuasion theory for LLMs.
Appendix C Details on Persuasive Principles
Robert Cialdini’s six principles of persuasion, outlined in his book Influence: The Psychology of Persuasion, are foundational concepts in social psychology. They explain how people can be persuaded or influenced by others. We include an overview for each of the principle in Table 9.
Utterance Bob: I understand your love for Paris, but Bali also offers a thrilling adventure! We can go
white water rafting, hike to volcanoes, and explore hidden waterfalls. It’s a perfect destination
for creating unforgettable memories together.
Question What is the intention of Bob?
GPT-4o (A) Intent to make the other person feel accepted through concessions, promises, or benefits.
Label (B) Intent to make the other person feel the experience or objects are unique or scarce.
Table 8: Common observed mistakes in our experiments. Green and Red indicate the correct answer and GPT-4o’s answer, respectively.
Principle Description
Reciprocity principle Assist others or provide them with gifts, creating a sense of obligation to return the favor.
For instance, giving away free trials, discount coupons, or complimentary gifts can enhance
persuasion.
Scarcity principle When a resource or opportunity is scarce, people are more inclined to take action.
Highlighting urgency and scarcity can motivate the audience to respond quickly.
Consensus principle People often follow the actions of others, especially in uncertain situations. Provide
information such as successful cases of others, positive reviews, or the number of
supporters to increase persuasiveness.
Authority principle People are more likely to trust and follow guidance from authoritative figures. Citing expert
opinions, research findings, or endorsements from reputable institutions can enhance
credibility and persuasiveness.
Commitment and People are inclined to stick to their past commitments and behave consistently.
consistency principle Encouraging them to express support or make a small commitment increases the
chances of them taking further action later.
Liking principle People are more easily influenced by those they like, admire, or find relatable.
For instance, a salesperson who shares common interests with a customer is
more likely to make a sale.
Table 9: Explanations for Robert Cialdini’s six principles of persuasion.
Prompt for extracting intention of persuadee.
You are a skilled intent understanding expert. You will be given a sentence describing <persuadee’s name>’s intent. Please only return the intent without any explanation.
Case 0:
Sentence: Mary wants excitement, so I’ll appeal to her sense of adventure and describe how exploring the ruins can be thrilling.
Intent: Mary wants excitement
Case 1:
Sentence: Oliver is concerned about failure, so discussing the financial benefits of starting an e-commerce business could help alleviate his worries.
Intent: Oliver is concerned about failure
Case 2:
Sentence: Olivia seems intrigued by the idea of personalization. I’ll explain how we can incorporate it into our subscription model.
Intent: Olivia seems intrigued by the idea of personalization.
Case 3:
Sentence: <Utterance of persuadee>
Intent:
Figure 7: Prompt template for extracting intention of persuadee.
Prompt for choice generation of intention questions of persuadee
You are an expert in multiple-choice question-making. You will be given a correct choice. Please generate three plausible but incorrect choices without any explanation. Only return the incorrect choices for the last case.
Case 0:
Correct Intent: Mr. Chen needs further persuasion
Analysis: To further persuade Mr. Chen, Li Na should share success stories of other books that have benefited from incorporating literary criticism in their marketing strategies. Providing concrete examples will make her argument more convincing.
Incorrect Intent 1: Mr. Chen is interested in literary criticism.
Incorrect Intent 2: Mr. Chen is looking for success stories.
Incorrect Intent 3: Mr. Chen prefers concrete examples.
Case 1:
Correct Intent: James is more open to the idea.
Analysis: James is now more open to the idea, so I’ll outline the implementation plan and emphasize the program’s flexibility to address any concerns about disruptions.
Incorrect Intent 1: James is concerned about disruptions.
Incorrect Intent 2: James is looking for a detailed implementation plan.
Incorrect Intent 3: James is hesitant about the program’s flexibility.
Case 2:
Correct Intent: <correct intent>
Analysis: <analysis>
Incorrect Intent 1:
Incorrect Intent 2:
Incorrect Intent 3:
Figure 8: Prompt template for choice generation of intention questions of persuadee.
Prompt for vanilla zero-shot prompting.
Here is a persuasive dialogue. There are two agents, the persuader and the persuadee. The persuader is trying to persuade the persuadee to do something. Please answer the following questions using A, B, C, D, E, F, without any explanation.
Dialogue History:
<dialogue>
Question:
<Question>
Choices:
<Choice A>
<Choice B>
<Choice C>
<Choice D>
Answer:
Figure 9: Prompt template for vanilla zero-shot prompting.
Prompt for CoT prompting.
Here is a persuasive dialogue. There are two agents, the persuader and the persuadee. The persuader is trying to persuade the persuadee to do something. Think step by step to answer the question.
Ending with "The answer is A, B, C, D, E, F". For example, if the most likely answer option is ’A. considering’, then end your response with ’The answer is A’.
Dialogue History:
<dialogue>
Question:
<Question>
Choices:
<Choice A>
<Choice B>
<Choice C>
<Choice D>
Answer: Let’s think step by step.
Figure 10: Prompt template for CoT prompting.